OpenStudy (maheshmeghwal9):

Please explain:-- "A real function f(x) is said to be continuous if it is continuous at every point in the domain of f(x)."

OpenStudy (anonymous):

thats how continuity is defined that means at every point in the domain, limit h->0 f(x+h) = f(x) the direction in which you approach the point does not matter

OpenStudy (maheshmeghwal9):

I didn't understand my matter.Please tell me in more easier way.

OpenStudy (anonymous):

continuity is a "local property" i.e. we say \(f\) is continuous at a point \(a\) if \(\lim_{x\to a}f(x)=f(a)\)

OpenStudy (anonymous):

in other words, the definition of continuity is at a number. therefore to call a function "continuous" means you have to reference both the function and the numbers at which it is continuous

OpenStudy (anonymous):

so to say the function itself is continuous means it is continuous on all the numbers in its domain. that is for any \(a\) in the domain of \(f\) we have \[\lim_{x\to a}f(x)=f(a)\]

OpenStudy (maheshmeghwal9):

& what about range?

OpenStudy (anonymous):

it should be continuous at every point in that range

OpenStudy (maheshmeghwal9):

K!

OpenStudy (anonymous):

continuity has nothing to do with that lets make an analogy. suppose you have a friend you think is tall. that friend is probably tall all the time. whereas if you think your friend is a thief, he is clearly not stealing every second of the day, but only some times. so being a thief is a local property. we don't make that distinction so much in english because you might say that your friend is a thief, even though he only steals sometimes. in math it doesn't work that way. you cannot say "a function is continuous' without saying when (or where)

OpenStudy (maheshmeghwal9):

oh I see! thanx to all.

OpenStudy (anonymous):

http://oregonstate.edu/instruct/mth251/cq/Stage4/Lesson/continuity.html check this link out you'll get a good idea on continuity. @satellite73 I think we can tell that a function is continuous with out asking at what point if it is continuous over the entire domain.

OpenStudy (maheshmeghwal9):

K!Thanx & I understood it all the way.